This problem is from a worksheet. Find all zeros of each function. If a zero has a multiplicity greater than on, please note the multiplicity.
f(x)= x^3+12^2-16
I have tried to solve using synthetic division, but cannot find a number that makes it zero, I have tried +-{1,2,4,8,16} and I put a zero in place where there is a missing degree...but still cannot find answer. My p/q is probably wrong, but I am pulling my hair out here trying to figure it out. Thanks for the help.

Answer by venugopalramana(3286)

This problem is from a worksheet. Find all zeros of each function. If a zero has a multiplicity greater than on, please note the multiplicity.
f(x)= x^3+12x^2-16...I THINK IT IS 12X^2..OK?
I have tried to solve using synthetic division, but cannot find a number that makes it zero, I have tried +-{1,2,4,8,16} and I put a zero in place where there is a missing degree...but still cannot find answer. My p/q is probably wrong, but I am pulling my hair out here trying to figure it out. Thanks for the help.
NOTHING WRONG WITH YOU.YOU ARE CORRECT.IT HAS A ROOT WHICH IS NOT RATIONAL.HENCE YOU ARE NOT GETTING THE ANSWER BY TRYING 1,2,4,8,16...
THE ROOT IS ABOUT 1.1 APPROXIMATELY.I AM NOT SURE WHICH METHODS YOU HAVE BEEN TAUGHT ,WHEN THE CUBIC HAS AN IRRATIONAL ROOT .JUST GIVE ME YOUR BACK GROUND AND I SHALL TRY TO HELP.ONE EASY WAY IS TO SOLVE GRPHICALY.THAT IS DRAW A GRAPH FOR Y=X^3 AND ANOTHER GRAPH FOR Y=16-12X^2 AND FIND THE POINT OF INTERSECTION.